The sum of two numbers is 84. The second is 9 more than 4 times the first. What are the two numbers?
Answer by stanbon(75887)

The sum of two numbers is 84. The second is 9 more than 4 times the first. What are the two numbers?
--------------
Let one of the numbers be "x".
The other one must be "84-x"
-------------------
EQUATION:
2nd is 4*1st + 9
84-x = 4*x+9
5x= 75
x=15 (1st number)
84-x=69 (2nd number)
--------------
Cheers,
Stan H.